OpenStudy (anonymous):

. The rate r at which a cricket chirps depends linearly on the temperature t. If the temperature is 50 degrees Fahrenheit, then the cricket chirps at the rate of 20 chirps per minute. Moreover, a rise in temperature of 10 degrees Fahrenheit causes the rate at which the cricket chirps to increase by 40 chirps per minute. Find a formula for r as a function of t.

OpenStudy (ybarrap):

slope = 10 degrees /50 chirps per min t = (10/50) *r + b when t = 50, r =20 so t - (1/5)r = b b = t - (1/5)r = 50 - (1/5)20 = 50 - 4 = 46 Ans. t = (1/5)r + 46